# Bridging the Social & Technical Divide in Augmentative and Alternative Communication (AAC) Applications for Autistic Adults

### v1

@@ 5.4.1. Trust
Nearly all of our participants (11) spoke about how using AAC depended on whether the environment is safe, what the comfort level is with the person they are speaking to, and social safety.
When I don’t use AAC is in some situations when there are people who have dramatic power over me who can drastically control my life, because my trauma tells me it is actively not safe to in those situations.
@@
Participants (2) don’t trust strangers to accept their AAC well; with good reason since they have been treated negatively in the past.
I didn’t use AAC last time I went to urgent care because I was worried the staff would treat me worse. […] I just used mouth words, and it went fine, but it felt like betraying my needs in a sense.
Other times it’s more nuanced. They feel pressured to be efficient in public (1) or are drained from matching others’ communication expectations (1).
Participants will often avoid social situations (4) rather than deal with strangers’ reactions to AAC. One participant (1) even avoids using AAC because they had abusive speech therapy when they were younger. This results in people not using AAC as much as they might need (2): “A lot of the time when I’m having to pull out AAC, I do the mental calculation of like ”do I really need to?””.
One participant (1), who mostly uses writing as their AAC, was the opposite. Because their speech is “unreliable”, they will primarily write with people who they are setting expectations with (1). Perhaps it is because of the lack of stigma against writing (vs an application) they are able to do this. Similarly, another participant (1) will either avoid social situations or use their notebook more.
@@ 5.4.2. Accommodating to current needs
There were themes of using energy to force speech or speak fluidly (4). Switching to AAC can feel like more (1) or less (1) energy than using mouth words. One participant (1) would occasionally fall into old habits of speaking orally by default. For another (1), who has an auditory processing disorder and mental fatigue, reading can be less tiring than listening. AAC can make it easier to overcome the “initiation cost” of figuring out when to talk (1).
@@
Participants (9) mentioned losing the ability to use mouth-words because of various internal or external stressors. Some (2) spoke about using AAC when they were overwhelmed or had a build up of stressors from throughout their day. Others (2) spoke about their co-morbid conditions that trigger the use of AAC. Some participants (2) use AAC depending on their energy and effort required.
It’s important for users to have access to AAC tools and be able to identify when their bodies are reaching a tipping point so that they know when to start using AAC and select what AAC to use (2). One participant (1) will tell their partner to hold themselves accountable, making sure they’re respecting their body’s boundaries.
5.4.3. Accommodating to the environment
Several participants (9) spoke about how the environment such as the weather and noise affects their AAC usage. Participants (9) spoke about using their AAC tool to show text in situations where the environment is loud (1), in a drive-thru (1), when they preferred not to use text-to-speech and when they have been misunderstood. Some participants spoke about the limitations of displaying text via AAC which include sun glare (1), font size, and concerns about privacy when handing one’s phone to read the text.
Some examples of situations where our participants might not have access to their device include: services where body is in a particular position and a device can’t be reached or looked at (e.g., salon (1), eye doctor (1)), when their hands are full already (2), while walking (1), when there’s no place to put a keyboard/larger device (2), wet environments such as showers and pools (2), and cooking (1).
5.4.4. Accommodating to the task
A quarter of our participants mentioned that they had encountered problems using AAC to give a formal presentation (2) or prefer a certain application for presenting (1). One participant had mentioned that passing around a piece of paper for a presentation is not practical, especially in a group presentation. However, they ran into technical difficulties when the AAC device they had at the time (DeltaTalker) spoke the entire presentation too fast to be heard. Another participant (1) complained about how Proloquo2Go’s text field only lets you see a sentence at a time, making it hard to edit an entire talk. Similarly, someone else (1) will use Proloquo4Text when they wants to use more technical vocabulary. These examples highlight an issue that AAC applications make an assumption that someone is using them for short sentences and not longer forms of communication.
Other users will switch to alternate ways to communicate that are appropriate for the task, beyond specialized AAC applications. One participant (1) has shifted to having most important work conversations via email. Another participant (1) will use printed-out laminated cards of pictures of groceries to help themself make shopping lists.
5.4.5. Accommodating to conversation partners
@@
Our participants would also switch AAC types depending on who they were communicating with. This need to switch can stem from accommodating the conversation partner’s disability, repairing if the partner cannot understand the TTS voice, or adapt to their partner’s skill of using a particular modality.
For instance, one participant (1) spoke about having options to adapt to others’ needs.
[Someone I know] is dyslexic and functionally illiterate. So needing to have both reading and auditory as options easily [available] and how to not have it overloading to switch between those for everyone based on what’s needed […] [It’s] a lot of: just what is needed in this moment? And having backups for backups and just being [ready] to switch in any moment.
Another participant (1) will switch to texting if their spouse doesn’t have their hearing aids in, or show them text if they don’t have their phone. A third participant (1) told of a story where they showed the text on their phone to a cashier, but she couldn’t see it and needed to take the phone to bring it close to her face. In order to communicate when their conversation partner can’t understand the TTS from their AAC, participants will change the speed of the TTS voice (1) or use chat/text to write out what was said (2).
AAC users will also change their conversation mode to adapt to someone’s skill level. One participant (1) is very comfortable with American Sign Language [^9] We recognize that ASL is a full language in its own right. However, it seems like our hearing participants use ASL more as a way of augmenting English., but they will write if people don’t sign and will practice signing if their conversation partner needs practice. Participants (2) also mentioned switching to whatever application their conversation partner is using in order to model/teach them how to use it.
@@ 5.5. Benefits
Of course people wouldn’t gravitate toward finding and using AAC if there weren’t significant benefits. The themes we found here include Organizing thoughts, Agency and identity, Self-accommodation, Community, and Fun.
@@ 5.5.1. Organizing thoughts
Participants (8) mentioned various ways in which AAC helped them “collect their thoughts”. The benefits can span from treating their AAC as an automatic note taking app (1) and keeping track of what they said (1) to getting around choice paralysis (1). A few participants will use their AAC to draft messages to be shared later via text or online (2).
Some autistic people have difficulty determining when it’s their turn in a conversation (Ochi et al., 2019). One participant (1) will use their AAC to type up what they would want to say, whether or not they will use it in a conversation. With having what they want to say out of the way, they can focus on when will be the right time to say it instead of keeping track of both what and when.
5.5.2. Agency and identity
Several participants (9) spoke about how AAC enabled them to communicate their needs and act independently. Using AAC has helped participants feel “like a whole person” (1) by letting them:
@@ 5.5.2. Agency and identity
  • Decide what they want to say and how (1)
  • Speak for themselves and make their own choices (3)
  • Talk about heavy, emotional things (1)
Sometimes all it takes is having the ability to say something when it’s hard to otherwise (5): “Just still being able to contribute even on a bad day.
AAC can also become closely intertwined with a user’s identity:
@@ 5.5.3. Self-accommodation
Since communication is so core to our experience as humans, some of our participants have structured their entire lives around accommodating their AAC use. One participant (1) has made it so that the majority of their “heavy lifting conversations” are via email, and another participant (1) has structured their life around AAC such that they “don’t run into the edges very often anymore”.
More than half of our participants (7) rarely even use mouth-words anymore, except with people they are close to (2), around people who don’t support its use (3), when they can’t access it (2), or when they need to communicate in a certain way such as quickly or with emotion (1).
Participants feel more confident speaking orally when they know that they have AAC as a backup (2). One (1) makes sure that they has their AAC with them when they’re in public. Another (1) will bring their AAC along in case the day becomes overwhelming.
5.5.4. Community
AAC gives users the ability to feel socially connected to others (3). Our participants also reported making friends with other AAC users (2) or other autistic people since they don’t judge others for using AAC (1). Our participants also feel a need to give back to their communities. A couple advocate for AAC use (2) or are an autism self-advocate at local medical schools (1). Participants reported helping out other disabled groups as well. One participant (1) had transcribed for deaf teens at an event in the past, and another participant (1) who is for all intents and purposes deafblind themself, is training to be a deafblind counselor.
5.5.5. Fun
A couple of participants even spoke to the fun aspects of AAC, playing with it and talk to pets (1) or even having AAC as a special interest [^10] In the autism community, a special interest is a topic or hobby that deeply fascinates someone, and they try to learn all they can about it. (1).
5.6. Access
Due to autistic adults finding and accessing AAC later in life, there is an additional overhead that autistic children do not run into—either because they are still a part of the US education system and therefore can more easily access speech-language pathologists, or their parents and caregivers take over the details regarding things like price and setup.
@@ 5.6.1. Institutional support
Support from speech-language pathologists (SLPs) or other institutions such as government programs in place to help people who use AAC or who are disabled in general was mixed. Several participants (4) spoke about how speech therapists/speech-language pathologists were unhelpful or even abusive, hindering their ability to use AAC or find the right AAC for themselves.
In my experience, […] speech therapists were not helpful. […] There’s a lot they can do to acclimate folks to AAC better. Understanding what it means to have part-time speech or unreliable speech, how it shows up, and how to work with it instead of against it—so that you’re working on total communication, rather than just speech because technically they’re speech and language therapists.
On the other hand, SLPs can create opportunities for AAC users. One participant (1) had an SLP that let them try out an expensive AAC so that they could decide if it was right for them or not first.
While one participant had found SLPs unhelpful, they spoke a lot about benefiting from state-level support. They were able to borrow technologies over the years that helped obtain AAC without having to buy it themselves. However, one participant (1) had run into people at a state-level disability organization who were not “taught how to treat us as humans”.
@@ 5.6.2. Pricing and inaccessible operating systems
Some participants spoke about being prevented from getting the AAC that they thought would best fit their needs because it was only available on an operating system that they did not have access to. They either wanted an application that was supported by multiple operating systems (i.e., cross-platform) (1) or was not being constrained to one operating system (3). This second point came up often in reference to AssistiveWare’s Proloquo suite.
Actually existing at all on Android. I feel like that has to be said actually. I actually trialed both Proloquo2Go and Proloquo4Text and found neither of them worked for me, even beyond the fact that iPads don’t generally work for me.
These applications are only developed for iOS. Therefore, users would need to invest in an Apple product, which is typically more expensive than an Android tablet or phone. One participant (1) felt discouraged from ever bringing their iPad around for fear of breaking it because it was so expensive. Some participants mentioned having a dedicated tablet—typically an iPad—for AAC (4), while others are interested in getting one eventually (1).
Similarly, even more participants (10) spoke explicitly about how affordability played an important role in choosing an AAC application. Participants would gravitate toward applications that are free (4), particularly if they are getting a separate app for their phone (1). There are downfalls of using a free app though; namely, fewer features (1). This limiting of features is especially seen in the text-to-speech voice options (3). Even applications that cost money to use will have “premium” voices that cost extra.
It’s also pretty messed up that they make you spend 40 more dollars, and you don’t even know for sure you’re going to like the voices in the expansion.
Other times, participants will pay for expensive apps when there is a big sale (as AssistiveWare apparently does annually; 2), are occasionally gifted them (2), or just don’t use the AAC they think will help them because it’s too expensive (1). Only very rarely when they are concerned about price will they buy it themselves (1). One participant would just get whatever AAC their mom could afford when they were a child.
One participant (1) ran into a weird intersection between operating system incompatibility, price, and government assistance. They were interested in trying Speech Assistant application, but “that’s free on Android and \$20 on iOS, and I’m not paying \$20 for software that’s free on another platform just because I was able to get the State to buy me the platform most people prefer for AAC”.
5.6.3. Universal access
Some participants (3) spoke about the need for AAC to be “more accessible to more people”. AAC should be integrated into society (1) and autistic adults should have access to SLP professionals (1). Additionally, we heard themes about a need for increased awareness regarding available AAC options, with one participant discussing lesser-known features that are already available to people:
And then the other thing […] about AAC that people don’t know is that a lot of smartphones have built-in features that they don’t know are there, like Text Relay […] for being able to make your own calls, and we need to help folks learn about those resources. […] Like, the government funds that, and people just don’t know it’s a thing.
Participants (2) discussed being introduced to AAC much later in life than they would have liked. Even after knowing about AAC and using it some participants (2) expressed uncertainty as to whether an AAC app existed with the features that they’re interested in.
5.6.4. Setup effort
Once people get access to an AAC application, they spend a lot of time setting the app up and making boards to fit their needs (4). With so many complicated pieces, it’s difficult for users to remember where everything is and what it can do (3). It makes for a big overhead for starting AAC or moving to another AAC application (2). One participant (1) looks at other apps or the boards of other users to get help making their own boards, but they wish that there was a companion app for their computer where they can more easily edit boards.
@@ 5.7. Stumbling Blocks for Continued Use
Once our participants found and setup their selected AAC applications, there were still major issues that they would encounter. These issues might discourage users from using particular features of the apps or might make them disregard the application entirely.
@@ 5.7.1. Speed
Nearly all participants (11) brought up the issue of speed with AAC, wanting “to be able to communicate faster”(1) and keep up in conversations (3). Even patient conversation partners can sometimes forget what an AAC user might be responding to (1) or get distracted by others in the conversation (1). We could see an inconsistency in speed with how long the interviews themselves took. There was more variance in the length of the interviews for participants who used AAC (see Figure 2), which could depend on how comfortable the participant was with their AAC and/or whether they scripted their answers to the questions ahead of time.
Figure 2. A box plot depicting the difference between interview times for participants who primarily used AAC during the interview (right, orange) and those who did not (left, blue). This data does not include the one participant who interviewed completely asynchronously.
[…]
Speaking participants (5 people) took on average 1.045 hours (standard deviation of 0.05) for the interview, and non-speaking participants (6 people) took on average 1.5197 hours (SD 0.6).
This issue of speed can stem from troubles with locating button words on the board (6), but there also just seems to be an order for how fast someone can communicate based on what method they’re using. Oral speech seems to be faster than any AAC (1) and can speed up AAC use (1), but in the hierarchy of AAC, external keyboards are faster than on-screen keyboards (3), swipe keyboards are faster than typing letter by letter (1), and spelling things out letter by letter—either on an on-screen keyboard (1) or a letter board (1)—is the slowest. One participant (1) also found hand writing to be faster than phone typing.
Although on-screen typing/text-based AAC might be used more often than symbol-based AAC for complex words or faster communication (1; see Section 5.1 for more discussion about symbol vs text input), others wish that they could type faster (2). Having an onscreen keyboard option for spelling out words to look up where a button might be or to avoid having to look for a button (3) or using auto-complete/predictive text can help with faster button input (3). The important characteristic here being that the keyboard can be easily accessed without significantly disrupting app use (1) or motor planning (1). Issues with finding buttons are exacerbated for people who only use AAC sometimes.
“I would need to use it regularly enough to memorize the symbol locations. Proloquo’s search feature, at least, does highlight the specific buttons to press to get to the requested word, which I think is really handy.”
There are also extra issues that can slow down communication, such as an app slowly loading (1), not having the board completely setup (1), or having to switch apps in the middle of a conversation (1).
@@ 5.7.2. Voice quality
Most participants (10) expressed their concerns with text-to-speech (TTS). Seven participants (7) felt that the available TTS voices were often hard to understand or hard to hear in general. They shared instances—such as during interactions with doctors (1) or in meetings (1)—in which these voices led to misunderstandings. Two participants (2) attempt to get around misunderstandings by changing the speed of the voice to be slower. Other misunderstandings can come from their conversation partners who are not used to synthesized voices (1) or have auditory processing disorders (APDs) (1). One participant accommodates to their partner with an APD by using a voice that has similar characteristics to the dialect of the area.
A couple participants felt that the quality of the voices was simply too poor to use them (2).
I’ve had times where I tried to pull out the AAC in a group conversation […] and then people hear it and laugh because they think I’m making a joke. […] It’s just like no one really knows how to react to it, especially when you’re a part time user. And so you’re like ‘sorry, I showed up more disabled today than you know me as’.
A few participants (3) wished their AAC could be heard over the phone. Another participant (1) has been hung up on in the past because people thought they were a robot. Some participants (4) get around this by using TTY-Based [^11] https://www.fcc.gov/tty-relay or IP Relay [^12] https://www.fcc.gov/ip-relay, a free service where state-certified employees read messages transmitted via texting or the internet, respectively. This gives people who would rather communicate via text a human voice to “speak” with over the phone. As far as we know, there are no AAC applications that integrate with Relay. One participant (1) explicitly wished their AAC could.
One participant (1) spoke about how the TTS lacks emotion, while another used a child’s voice because it was more expressive (1). Additionally, the voices from the device can get masked by background noise because they simply can’t become loud enough (4). Participants who primarily show text consider TTS as a bonus (1).
@@
[…]
The UbiDuo device has a slightly smaller than normal sized keyboard with a small (3-5 inches wide) backlit screen above it. It looks like the screen can fold down toward the keyboard, but whether it’s folded or not, the base looks like it might be at least an inch thick.

### v3

@@ 5.4.1. Trust
Because autism is an invisible disability and some people have the ability to pass as neurotypical, using AAC can “out them” as autistic. Nearly all of our participants (11) spoke about how using AAC depended on whether the environment is safe, what the comfort level is with the person they are speaking to, and social safety.
I don’t use AAC […] when there are people who have dramatic power over me who can drastically control my life, because my trauma tells me it is actively not safe to in those situations.
@@
Other participants do not trust strangers to accept their AAC well due to being treated poorly in the past (3), and can feel pressured to match others’ communication expectations (2). Donaldson et al. (2021) also found trust to be a major theme influencing autistic adults’ choice of AAC. As a result, participants (4) will often avoid social situations rather than deal with strangers’ reactions.
@@ 5.4.2. Accommodating to current needs
Having a dynamic disability means that one’s needs are not constant throughout their life or even throughout their day. Like any accessibility tool (Shinohara and Wobbrock, 2011), autistic adults need to find the right AAC for their current needs (Donaldson et al., 2021). Here, there were themes of using energy to force speech or speak fluidly (4). Switching to AAC can feel like more (1) or less (1) energy than using mouth words. One participant (1) would occasionally fall into old habits of speaking orally by default. AAC can make it easier to overcome the “initiation cost” of figuring out when to talk (1).
@@
Zisk and Dalton (2019) note that intermittent speech–losing speech due to sensory stressors or co-occurring conditions–is rarely researched. Our participants (9) mentioned losing the ability to use mouth-words because of various internal or external stressors. Some (2) spoke about using AAC when they were overwhelmed or had a build up of stressors from throughout their day. Others (2) spoke about their co-morbid conditions that trigger the use of AAC. Some participants (2) use AAC depending on their energy and the effort required. It is important for users to have access to AAC tools and be able to identify when their bodies are reaching a tipping point so that they know when to start using AAC and select what AAC to use (2). One participant (1) will ask their partner to help them hold themselves accountable so that they respect their body’s boundaries.
5.4.3. Other conversational contexts
Donaldson et al. (2021) have seen that autistic adults’ AAC selection depends on the conversational context. Our interviews back this up as well:
Accommodating to the environment.
Several participants (9) spoke about how the environment such as the weather or noise levels affects their AAC usage. Participants (9) spoke about using their AAC tool to show text in situations where the environment prohibits TTS from being heard. There are also situations in which participants do not have access to their device because of what they were doing (5), there was no room (2), or they were concerned that the device would be damaged – e.g., around water (2).
Accommodating to the task.
Like other AAC users have found (Obiorah et al., 2021), AAC is not always practical to use for every task. A quarter of our participants (3) mentioned constraints with using AAC to give presentations to groups of people, where they would run into clunky interfaces or bugs. Other users (2) will switch to alternate ways to communicate that are appropriate for the task beyond specialized or high-tech AAC applications, such as using email for important work communications or laminated cards for making grocery lists.
Accommodating to conversation partners.
@@ 5.4.3. Other conversational contexts
Our participants would also switch AAC types depending on who they were communicating with. This need to switch can stem from accommodating the conversation partner’s disability (3), repairing if the partner cannot understand the TTS voice (3), or adapting to their partner’s skill of using a particular modality or application (2).
@@ 5.5. Benefits
Of course adults would not seek out AAC if there were not significant benefits. AAC gives users the ability to feel socially connected to others (3), with our participants making friends with other AAC users (2). One participant even stated having AAC as a special interest [^10] In the autism community, a special interest is a topic or hobby that deeply fascinates someone, and they try to learn all they can about it.. In this section we will talk about three main categories where people found AAC helped them beyond using it as a communication tool: Organizing thoughts5.5.1), Identity5.5.2), and Self-accommodation5.5.3).
@@ 5.5.1. Organizing thoughts
Participants (8) mentioned various ways in which AAC helped them “collect their thoughts”. Their AAC can be used to reference what they said in a previous conversation (1) or earlier in the current conversation (1), to overcome choice paralysis (1), or to draft messages to be shared later via text or online (2).
Some autistic people have difficulty determining when it is their turn in a conversation (Ochi et al., 2019). One participant uses their AAC to type up what they would say, whether or not they will use it in a conversation. With having what they want to say out of the way, they can focus on when to say it.
5.5.2. Identity
Several of our participants (9) spoke about how AAC enabled them to communicate their needs and act independently. Using AAC has helped participants feel “like a whole person” by letting them:
@@ 5.5.2. Identity
  • Speak for themselves (3)
  • Talk about emotional things or in emotional situations (3)
  • Have the ability to communicate when it would otherwise be difficult (5) (“Just still being able to contribute even on a bad day.”)
Subsequently, AAC use can become closely intertwined with a user’s sense of self:
@@ 5.5.3. Self-accommodation
Since communication is so core to our experience as humans, some of our participants have structured their entire lives around accommodating their AAC use. One participant (1) has made it so that the majority of their “heavy lifting conversations” are via email, and another participant (1) has structured their life around AAC such that they “don’t run into the edges very often anymore”. More than half of our participants (7) rarely even use mouth-words since starting AAC, except with people they are close to (2), around people who do not support its use (3), when they cannot access it (2), or when they need to communicate in a certain way, such as quickly or with emotion (1). Participants also feel more confident speaking orally when they know that they have AAC as a backup (2).
5.6. Access as an Adult
Due to many autistic adults finding and accessing AAC later in life, there is an additional overhead that autistic children do not run into—either because they are still a part of the US education system and therefore can more easily access speech-language pathologists (Institutional support, §5.6.1), or because their parents and caregivers take over the details regarding things like setup (§5.6.2) and pricing (§5.6.3).
@@ 5.6.1. Institutional support
Opinions about support from speech-language pathologists (SLPs) or other institutions—such as government programs in place to help people who use AAC or who are disabled in general—were mixed. Some participants (4) spoke about how speech therapists or SLPs were unhelpful or even abusive, hindering their ability to use AAC or find the right AAC for themselves.
In my experience, […] speech therapists were not helpful. […] There’s a lot they can do to acclimate folks to AAC better. Understanding what it means to have part-time speech or unreliable speech, how it shows up, and how to work with it instead of against it—so that you’re working on total communication.
On the other hand, SLPs can create opportunities for AAC users. One participant had an SLP that let them try out an expensive AAC so that they could decide if it was right for them or not first. Another person (1) wished that there was more access to speech-language pathologists as an adult so that they might get more help for their AAC use, a barrier also seen by Zisk and Dalton (2019). Furthermore, several participants (5) expressed concerns that could be addressed with an SLP, such as whether an AAC app exists with the features that they like, how to get started with a particular type of AAC, or figuring out if they like certain features.
Regardless, some participants (3) spoke about the need for AAC to be “more accessible to more people”, which can potentially be addressed via institutional supports. A third participant (1) had found SLPs unhelpful, but they spoke a lot about benefiting from state-level support; they were able to obtain AAC technologies over the years without having to buy them themself.
5.6.2. Setup effort
Because many of our participants were introduced to AAC as an adult outside of institutional support, they had to find and setup their applications themselves. Once people get access to an AAC application, they spend a lot of time setting the app up and making boards to fit their needs (4). With so many complicated pieces, it is difficult for users to remember where everything is and what it can do (3). It makes for a big overhead for starting AAC or moving to another AAC application (2).
@@ 5.6.3. Pricing and inaccessible operating systems
Some participants (4) spoke about being prevented from getting the AAC that they thought would best fit their needs because it was only available on an operating system that they did not have access to. AssistiveWare’s Proloquo suite of applications are only developed for iOS. Therefore, users would need to invest in an Apple product, which are typically more expensive than Android devices: “I don’t take my iPad with me […] most of the time. It was expensive, and I don’t want to break it.
Similarly, most participants (10) spoke explicitly about how affordability played an important role in choosing an AAC application, a pain point echoed by Zisk and Dalton (2019). Participants will avoid paying full price for expensive applications if possible (5) and would gravitate toward applications that are free (4). Applications will occasionally charge additionally for text-to-speech voice options (3).
@@ 5.7. Stumbling Blocks for Continued Use
Once our participants found and setup their selected AAC applications, there were still major issues that they would encounter. These issues might discourage users from using particular features of the apps or might make them disregard the application entirely. Most of these concerns, with the exception of External support5.7.1), do not appear to be autistic-adult–specific and could be shared with other AAC-using populations. The other, more common stumbling blocks include Speed5.7.2), Voice quality5.7.3), Limitations of showing text5.7.4), Device restrictions5.7.5), and Unreliability5.7.6).
5.7.1. External support
Because our participants mostly lacked institutional support, customer support can be key to whether or not a user will continue using an application. Participants who experienced good customer support (2) felt that they can rely on the company to continually improve the app. On the other hand, participants who experienced poor customer support (3) will switch applications, if possible, or suffer in silence since their concerns are not being actively addressed by the developers.
@@ 5.7.2. Speed
Speed has long been a concern in AAC use (Newell et al., 1998). Like other adult AAC users (McKillop, 2018; Dai et al., 2022), nearly all participants (11) mentioned the issue of speed with AAC, wanting “to be able to communicate faster” and keep up in conversations (3). Even patient conversation partners are not immune to distractions (2) and can sometimes forget what an AAC user might be responding to or get distracted by others in the conversation. This issue of communication speed can stem from issues with locating button words on the board (6) or the input method that is being used (5).
@@ 5.7.3. Voice quality
Most participants (10) expressed their concerns with poor text-to-speech (TTS) quality, which is a common issue with AAC users (McKillop, 2018). Seven participants (7) felt that the available TTS voices were often hard to understand or hard to hear in general. Two participants (2) attempt to get around misunderstandings by changing the speed of the voice to be slower. Other misunderstandings can come from when their conversation partners are unable to process a synthesized voice (2)
A couple participants felt that the quality of the voices was simply too poor to use them (2) or were concerned about the ability to be expressive with the voice (2). Other TTS restrictions include not being able to use AAC to talk to people over the phone (3) or to play the voices loud enough (4).
Figure 3 is a picture of a participant holding their UbiDuo device. The device has a slightly smaller than average sized keyboard with a small backlit screen above it. The screen can fold down, but it looks like it wouldn’t help it be any thinner.